a. The results/ observation of several tensile tests performed. Compare and contrast theoretical material properties: modules of elasticity (205GPa), yield stress (300MPa) and ultimate tensile stress (500MPa) of metal (ferrite stainless steels Grade 409) of the specimens 1 and specimen 16 and using the destructive tensile test results in the graph for relevant specimens.

b. Describe how the Archimedes’ principle-based test can be used to determine the density of two material samples using the following data. Compare your results with the given densities if there are variations, and mention possible reasons for that.

Proposed an NDT method can be used to prove your reasons. Compare and contrast your method with the method above.

A. Metal sample 1 data: Steel cylindrical sample (diameter 20mm, length 100mm and mass 252.80grams, density 8050kg/m3)

B. Non-metal sample 2 data: Nylon hollow shaft (internal diameter 10mm, outer diameter 16mm, length 300mm, mass 42.40g and density 1.15g/cm3)


1
Expert's answer
2021-07-19T03:14:48-0400

a.

Theoretical given

"E= 205GPa\\
S_{yt}= 300 MPa\\
S_{ut}= 500 MPa\\"

Practical elasticity

Modulus of elasticity =Slope of the curve

"\implies \frac{350-0}{\frac{0.03}{100}-0}\\
F= \frac{11666.66}{10^{-2}}= 1166.66 GPa"

Yield stress is 350 MPa and ultimate stress is 475 MPa from the graph

Specimen 1

Theoretical given

"E= 205GPa\\
S_{yt}= 300 MPa\\
S_{ut}= 500 MPa\\"

Practical results

The slope is the same for all curves, "E=1166.66 GPa"

Yield stress is 350 MPa and ultimate stress is 400 MPa from the graph


b.

The density of an object can be found out by using the Archimedes principle. The buoyant force on a submerged object is equal to the weight of the water displaced. 

The apparent weight can be found out from "W_{app}= \rho Vg" where "\rho" is the density, v is the volume of the object and g is the acceleration due to gravity. 

By knowing apparent weight we can find the density of the object. Therefore we have to find the apparent weight to find the density of object displaced.

"W_{app}= (\rho_s-\rho_l) Vg\\
 W_{app}= \frac{Mass}{\rho_s*V }"